OpenStudy (anonymous):

How many yards of material from a 24 yard length of cloth remain after 3 pieces, each 3 1/2 yards long, and 5 pieces each 2 1/4 yards long are removed?

OpenStudy (anonymous):

Length remaining = total length - removed length length = 24 - (3*3.5 + 5*2.25) = 24 - (10.5 + 11.25) = 24 - 22 = 2 yards